nvest in Project 1 because it has the higher IRI. Invest in Project 2 because it has the higher NPV 8 I projects below. The cash flows as well as the NPV and IRR t 13 Consider the two projects below 1or the two projec 10 percent. ts are given. For both projects, the required rate of return is Cash Flows Year Project 1 Project 2 4 36 0175 NPV 14.12 19.53 IRR (%) 16.37 15.02 1 2 3 100 36 36 36 -100 0 What discount rate would result in the same NPV for both projects? A A rate between 0.00 percent and 10.00 percent. B A rate between 10.00 percent and 15.02 percent. C A rate between 15.02 percent and 16.37 percent.
Answer #1

Option b is correct

NPV A = NPV B

NPV= C0+ CF1/(1+r)^1 + CF2/(1+r)^2 …………CFn/(1+r)^n

Hence

-100 + 36/(1+r)^1 + 36/(1+r)^2+ 36/(1+r)^3 + 36/(1+r)^4 = -100+ 175/(1+r)^4

36/(1+r)^1 + 36/(1+r)^2+ 36/(1+r)^3 + 36/(1+r)^4 = 175/(1+r)^4

This implies:

PMT = -36, N=4, FV = 175 and we have to find the Rate.

Find I/Y in financial calculator

As 13.16%